# TransKD: Transformer Knowledge Distillation for Efficient Semantic Segmentation

### v3

@@ III-A Overview
Existing semantic segmentation transformers generally follow either an isotropic structure [7, 69, 70] or a four-stage hierarchical structure [25, 26, 34, 50, 71]. Since the latter one is better at modeling multi-scale long-range information [24, 50], it is more versatile in dense prediction tasks such as pixel-wise semantic segmentation. Thereby, in this work, knowledge distillation is carried out on four-stage transformers (see Fig. 4).
@@
This “patchification” is the key ingredient of vision transformers as it allows us to cast vision-based tasks as sequence-to-sequence problems for which this type of models was initially built [14]. The patch embedding module splits the input images or feature maps $\mathbf{I}\in{\mathbb{R}}^{C{\times}H{\times}W}$ into patches and transforms these patches (mostly via a convolution operation) into a representative patch embedding $\mathbf{E}\in\mathbb{R}^{N\times C^{\prime}}$, where $N$ is the number of patches and $C^{\prime}$ the number of the channels. As often highlighted in previous research, the resulting patch embeddings convey important information about long-range sequential relationships [7, 72], location priors [7, 13, 69], and local continuity cues [25, 26, 71].
@@
where the feature map $\mathbf{F}\in\mathbb{R}^{{C^{\prime}\times{H^{\prime}\times{W^{\prime}}}}}$ marks the output of the current transformer stage and therefore the input to the next layer. A positive consequence of self-attention exploring all the patch tokens with no restriction to a certain receptive field is that the global context also being covered in the resulting feature map. Distilling knowledge from both, feature maps and patch embeddings, allows us to track complementary information within the transformer, such as spatial and sequential relationships.
Motivated by this, we propose Transformer-based Knowledge Distillation (TransKD) – a novel transformer-to-transformer knowledge distillation framework with the overall goal of building small yet accurate semantic segmentation transformers less reliant on time-consuming pre-training. TransKD centers around two types of knowledge transfer: (1) patch embedding distillation and (2) feature map distillation. The patch embedding distillation is performed at each of the four transformer stages while the feature map distillation leverages the relation-based knowledge distillation to gather information from within-stage and cross-stage feature maps. An overview of this core pipeline is depicted in Fig. 4(a).
@@
where $L_{embd}^{m}$ and $L_{fm}^{m}$ refer to the patch embedding loss and the feature map loss between the $m$-th stage of the teacher and the student. As mentioned before, we explore the typical four-stage transformer in this work, i.e., $M=4$ in our case. $\alpha_{m}$ and $\beta_{m}$ are the $m$-th elements of $\mathbf{\alpha}$ and $\mathbf{\beta}$, which control the weights of the patch embedding loss and the feature map loss at each stage, respectively. $\mathbf{\alpha}$ and $\mathbf{\beta}$ are set as $\left[0.1,0.1,0.5,1\right]$ and $\left[1,1,1,1\right]$. In the following method descriptions, matrices with superscripts $in$ and $out$ refer to inputs and outputs of the proposed modules, and matrices with the superscript $mid$ refer to processed matrices within the modules.
In Sec. III-B, we present the design of our patch embedding distillation scheme, and in Sec. III-C we elaborate our feature map distillation method. Finally, in Sec. III-D, we describe the configurations of different variants of the TransKD framework.
@@ III-B Patch Embedding Distillation
To maximize the potential of transformer-to-transformer knowledge distillation, we consider that transformer-specific patch embeddings should be fully exploited. The patch embedding channels of the student and the teacher models are different and a dimensionality transformation is an essential prerequisite step for patch embedding distillation at different network depths. To meet this requirement, we propose the Patch Embedding Alignment (PEA) module, which acts as a fundamental module for the channel dimension transformation. Additionally, we design the Global-Local Context Mixer (GL-Mixer) and the Embedding Assistant (EA) to optimize patch embedding distillation.
Patch embedding alignment. We conduct the Patch Embedding Alignment (PEA) through a linear projection for dimensionality transformation, after which the MSE loss between the student and teacher patch embeddings is calculated, as depicted in Eq. (4):
@@
where $\bm{E}_{m}^{S}$ and $\bm{E}_{m}^{T}$ are the patch embeddings at the $m$-th student and teacher stages. A learnable matrix $\bm{W}_{e}$ is applied to counter the dimensionality mismatch. PEA bridges the teacher and student transformers and facilitates multi-stage patch embedding streams, while only adding marginal computational complexity.
@@
Global-local context mixer. Patch embeddings deliver heterogeneous patterns [13, 26]. Therefore, the global context cues and the fine-grained local features, both of which are essential for semantic segmentation, cannot be fully distilled by using linear projection alone. To our knowledge, we are the first to introduce the two-branch design to the field of knowledge distillation in semantic segmentation. The global context is extracted via multi-head attention (left branch in Fig. 5) and the local information is extracted via a convolution layer (right branch in Fig. 5). Another convolution operation followed with sigmoid function $\sigma\left(\cdot\right)$ is used to control the information flow. Given an input patch embedding $\mathbf{E}_{m}^{in}$, the two-branch processing of global and local cues inside the GL-Mixer is formalized in Eq. (5):
@@
where $\mathbf{W}$, $\mathbf{V}$, $\mathbf{a}$, and $\mathbf{b}$ refer to learnable parameters for the convolutional operations. As shown in Fig. 5, the convolution operations are implemented via $3{\times}3$ convolutions and the entire GL-Mixer is lightweight. This way, given a representative patch embedding within a dense prediction transformer, both the semantically rich global context and the detail-rich local features can be harvested for distillation.
Figure 6: Combination of Embedding Assistant (EA) and Patch Embedding Alignment (PEA) in Fig. 4(c). The knowledge distillation pipeline through the EA modules and the student transformer blocks can be regarded as a medium-sized pseudo assistant model to provide supplementary distillation streams. $\mathbf{F}_{m}^{in}$ denotes the resulting feature map of the $m$-th stage, which is processed by the student’s embedding module to output the patch embedding of the next stage $\mathbf{E}_{m+1}^{in}$.
Embedding assistant. Knowledge distillation runs into problems when the size gap between student and teacher is significant [28, 29]. If the teacher becomes very complex, the student eventually lacks the sufficient capacity or mechanics to mimic its behavior despite receiving multi-stage hints. Mirzadeh et al. [28] introduce a Teacher Assistant (TA) of an intermediate size to bridge the size gap. TA acts as a mediator between the teacher and the student: it is distilled from the teacher model and then passes the knowledge to the student. TA follows the same structure as the student and the teacher, except for the numbers of hidden states and transformer layers [68]. The number of hidden states is adopted from the student, while the amount of the transformer layers corresponds to the one of the teacher. However, one weak spot of the TA paradigm is its higher computational cost and an additional training stage required for training the intermediate TA model.
@@ III-C Feature Map Distillation
Many knowledge distillation methods deploy a single-stage feature-based scheme for semantic segmentation [9, 18, 19], i.e., the information is exchanged at a single network depth. Our next consideration is that cross-stage relation-based knowledge distillation is important for efficient segmentation transformers, as it enables information transfer across different network depths. Similar ideas have shown success in the past, for example, in the case of hierarchical dense prediction transformers which utilize cross-stage feature maps [26, 50].
@@
We revisit Knowledge Review [10], which studies the connections across different levels of the teacher and the student CNN models. However, the Attention Based Fusion (ABF), which is the original feature map fusion module in Knowledge Review, dynamically aggregates feature maps via a spatial attention map, yet, overlooks the channel dimension within the attention maps. We therefore aim to design a complementary channel selection module. Our design is motivated by the following two aspects: 1) The channel-wise cues of the feature maps provide informative priors for feature extractors, and numerous previous works [73, 74] on feature map fusion, seek to fuse spatial and channel information via channel attention. Furthermore, Shu et al. [9] indicate that minimizing the channel-wise Kullback-Leibler (KL) divergence of the feature- and prediction maps is more efficient than using its spatial counterpart. 2) As the patch embedding distillation tracks dependencies among distant spatial positions via sequence-wise learning, the feature map distillation should engage more spatially holistic knowledge in channels. It is therefore important to highlight channels from certain stages via reweighting and account for cross-stage channel interdependencies. To achieve this, we propose the Cross Selective Fusion (CSF) module to adaptively fuse the information across stages via channel attention.
@@
Cross selective fusion. In this module, the feature maps from the current student stage and the next CSF are first resized to a uniform dimensionality and then fused via channel-wise attention. As shown previously in [73], feature extractions with different kernel sizes or modalities could be fused in an adaptive way. Similarly, we propose to fuse cross-stage feature maps for a relation-based distillation. As shown in Fig. 7, we first conduct a $1{\times}1$ convolution to implement the channel-wise transformation of the input feature map $\mathbf{F}_{in}^{m}$ and conduct interpolation to implement the spatial transformation of the feature map $\mathbf{F}_{mid}^{m+1}$ from CSF of the next stage. Their height and width are the same as the feature map of the $m$-th stage $\mathbf{F}_{in}^{m}$. The resulting ${\widetilde{\mathbf{F}}^{in}}_{m}$ and ${\widetilde{\mathbf{F}}^{mid}}_{m+1}$ are element-wise summed, as depicted in Eq. (6):
@@
where $\mathbf{A},\mathbf{B}\in{\mathbb{R}}^{C\times d}$ are two learnable matrices and $\mathbf{A}_{c},\mathbf{B}_{c}\in{\mathbb{R}}^{1\times d}$ are their $c$-th elements. $a_{c}$, $b_{c}$, and $\mathbf{F}_{mid,c}^{m}$ are the $c$-th elements of $\bm{a}$, $\bm{b}$, and $\mathbf{F}_{mid}^{m}$. $\bm{a}$ and $\bm{b}$ denote the soft attention vector for ${\widetilde{\mathbf{F}}^{in}}_{m}$ and ${\widetilde{\mathbf{F}}^{mid}}_{m+1}$, respectively. Further, we conduct a $3{\times}3$ convolution to convert $\mathbf{F}_{mid}^{m}$ to $\mathbf{F}_{out}^{m}$, which has the same dimension as the teacher’s feature map.
For training, we compute the HCL loss [10] between the outputs of the CSF and the teacher transformer block at each stage, as shown in Fig. 4.
@@ III-D Configuration
The core of our TransKD framework involves two knowledge distillation strategies: the patch embedding distillation (in Sec. III-B) and the feature map distillation (in Sec. III-C). To realize this design, we introduce four modules: PEA, GL-Mixer, and EA modules guide the patch embedding distillation, while the CSF feature map fusion module enhances the cross-stage relation-based feature map distillation.
Overall, CSF and PEA can be viewed as the fundamental components of TransKD. GL-Mixer operates on the deepest representative embedding, (i.e., the final fourth stage) producing highly heterogeneous patterns [26]. It reinforces the patch embedding distillation at the final stage by harvesting global and local context information in two branches. The EA modules are placed at each stage and serve as an embedding method to provide supplementary distillation streams by using an identical embedding method to the teacher’s and the student’s, with the teacher’s number of channels. It should be taken into account, that as different transformer architectures have different embedding methods, the framework leveraging EA does not yield full plug-and-play flexibility.
We implement and examine different variants of our proposed framework. Note, that we cannot isolate and verify the efficacy of each module when all the modules are combined together, as an integration of these components means many constraints and could lead to overfitting. We, therefore, focus on three configurations of the feature map distillation and the patch embedding distillation provided in Tab. I, which we refer to as TransKD-Base, TransKD-GL, and TransKD-EA. The structure of these three variants is provided in Fig. 4.
Table I: Configuration of TransKD variants.
@@ IV-A Datasets and Metrics
where $k$ is the number of classes, $p_{ij}$ is the number of pixels belonging to the class $i$ and classified as class $j$. We use MMCV [^1] MMCV: https://github.com/open-mmlab/mmcv to compute FLOPs and the number of parameters of our models.

### v4

@@ III-A Overview
Existing semantic segmentation transformers generally follow either an isotropic structure [11, 69, 70] or a four-stage hierarchical structure [25, 26, 34, 50, 71]. Since the latter one is better at modeling multi-scale long-range information [24, 50], it proves more versatile in dense prediction tasks such as pixel-wise semantic segmentation. Thereby, in this work, knowledge distillation is carried out on four-stage transformers, as illustrated in Fig. 4.
@@
This “patchification” is the key ingredient of vision transformers as it allows us to cast vision-based tasks as sequence-to-sequence problems for which this type of models was initially built [14]. The patch embedding module splits the input images ohes. Here, $H$, $W$, and $C$ represent the height, the width, and the number of channels of the input image or feature map of the embedding module, respectively. The module then transforms these patches (mostly via a convolution operation) into a sequence of representative patch embeddings $\mathbf{E}\in\mathbb{R}^{N\times C^{\prime}}$, where $N$ is the number of patches and $C^{\prime}$ the number of the channels. As often highlighted in previous research, the resulting patch embeddings convey important information about long-range sequential relationships [11, 72], location priors [11, 13, 69], and local continuity cues [25, 26, 71]. Both global context and local information contained in patch embeddings are relevant to semantic segmentation.
@@
where the feature map $\mathbf{F}\in\mathbb{R}^{{C^{\prime}\times{H^{\prime}\times{W^{\prime}}}}}$ marks the output of the current transformer stage and therefore the input to the next layer. A positive consequence of self-attention exploring all the patch tokens with no restriction to a certain receptive field is the global context also being covered in the resulting feature map. Distilling knowledge from both, feature maps and patch embeddings, allows us to track complementary information within the transformer, such as spatial and sequential relationships.
Motivated by this, we propose Transformer-based Knowledge Distillation (TransKD) – a novel transformer-to-transformer knowledge distillation framework with the overall goal of building small yet accurate semantic segmentation transformers less reliant on time-consuming pre-training. TransKD centers around two types of knowledge transfer: (1) patch embedding distillation and (2) feature map distillation. The patch embedding distillation is performed at each of the four transformer stages, while the feature map distillation leverages the relation-based knowledge distillation to gather information from both within-stage and cross-stage feature maps. An overview of this core pipeline is depicted in Fig. 4(a).
@@
where $L_{embd}^{m}$ and $L_{fm}^{m}$ refer to the patch embedding loss and the feature map loss between the $m$-th stage of the teacher and the student. As previously mentioned, we explore the typical four-stage transformer in this work, i.e., $M=4$ in our case. The variables $\alpha_{m}$ and $\beta_{m}$ are the $m$-th elements of $\mathbf{\alpha}$ and $\mathbf{\beta}$, which control the weights of the patch embedding loss and the feature map loss at each stage, respectively. The vectors $\mathbf{\alpha}$ and $\mathbf{\beta}$ are set as $\left[0.1,0.1,0.5,1\right]$ and $\left[1,1,1,1\right]$. In the following method descriptions, matrices with superscripts $in$ and $out$ refer to inputs and outputs of the proposed modules, while those with the superscript $mid$ refer to matrices processed within the modules.
In Sec. III-B, we present the design of our patch embedding distillation scheme, and in Sec. III-C, we elaborate our feature map distillation method. Finally, in Sec. III-D, we describe the configurations of different variants of the TransKD framework.
@@ III-B Patch Embedding Distillation
As previously discussed, the patch embedding incorporates positional information that enhances the feature maps. To maximize the potential of transformer-to-transformer knowledge distillation, we consider that transformer-specific patch embeddings should be fully exploited. Despite the specialized designs, the cumbersome teacher models consistently possess more transformer blocks for feature extraction and additional channels for knowledge retention. To align dimensions between student and teacher patch embeddings for loss function calculation and to introduce trainable parameters for learning channel-wise knowledge, we propose a fundamental module, Patch Embedding Alignment (PEA), to facilitate the simplest patch embedding distillation. Additionally, we design the Global-Local Context Mixer (GL-Mixer) and the Embedding Assistant (EA) to optimize patch embedding distillation.
Patch embedding alignment. We conduct the Patch Embedding Alignment (PEA) through a linear projection for dimensionality transformation. Subsequently, the Mean Squared Error (MSE) loss between the student and teacher patch embeddings is calculated, as depicted in Eq. (4):
@@
where $\bm{E}_{m}^{S}$ and $\bm{E}_{m}^{T}$ are the sequence of patch embeddings at the $m$-th student and teacher stages. A learnable matrix $\bm{W}_{e}\in{\mathbb{R}}^{C^{S}_{m}{\times}C^{T}_{m}}$ is applied to counter the dimensionality mismatch and transfer channel-wise knowledge from teacher patch embeddings to the student. $C^{S}_{m}$ and $C^{T}_{m}$ denote the number of channels in teacher and student models, respectively, at $m$-th stage. PEA bridges the teacher and student transformers and facilitates multi-stage patch embedding streams, while adding only marginal computational complexity.
@@
Global-local context mixer. Patch embeddings deliver heterogeneous patterns [13, 26]. Therefore, the global context cues and the fine-grained local features, both essential for semantic segmentation, cannot be fully distilled using linear projection alone. To our knowledge, we are the first to introduce the two-branch design to the field of knowledge distillation in semantic segmentation. The global context is extracted via multi-head attention (left branch in Fig. 5), and the local information is extracted via a convolution layer (right branch in Fig. 5). Another convolution operation, followed with sigmoid function $\sigma\left(\cdot\right)$, is used to control the information flow. Given the sequence of patch embedding at the $m$-th stage $\mathbf{E}_{m}^{in}$, the two-branch processing of global and local cues inside the GL-Mixer is formalized in Eq. (5):
@@
where $\mathbf{W}$, $\mathbf{V}$, $\mathbf{a}$, and $\mathbf{b}$ refer to learnable parameters for the convolution operations. As shown in Fig. 5, the convolution operations are implemented via $3{\times}3$ convolutions, and the entire GL-Mixer is designed to be lightweight. This way, given a sequence of representative patch embedding within a dense prediction transformer, both the semantically rich global context and the detail-rich local features can be harvested for distillation.
Figure 6: Combination of Embedding Assistant (EA) and Patch Embedding Alignment (PEA) in Fig. 4(c). The knowledge distillation pipeline through the EA modules and the student transformer blocks can be regarded as a medium-sized pseudo assistant model, providing supplementary distillation streams. $\mathbf{F}_{m}^{in}$ denotes the resulting feature map of the $m$-th stage, which is processed by the student’s embedding module to output the patch embedding of the next stage $\mathbf{E}_{m+1}^{in}$.
Embedding assistant. Knowledge distillation runs into problems when the size gap between student and teacher is significant [28, 29]. If the teacher becomes very complex, the student eventually lacks the sufficient capacity or mechanics to mimic its behavior despite receiving multi-stage hints. Mirzadeh et al. [28] introduce a Teacher Assistant (TA) of an intermediate size to bridge the size gap. TA acts as a mediator between the teacher and the student: it is distilled from the teacher model and then passes the knowledge to the student. TA follows the same structure as the student and the teacher, except for the numbers of hidden states and transformer layers [68]. The number of hidden states is adopted from the student, while the amount of the transformer layers corresponds to the one of the teacher. However, implementing Teacher Assistant for knowledge distillation is a two-step process. An additional assistant model TA should be trained, which is larger than the student model, thus introducing extra computation overhead and learning time.
@@ III-C Feature Map Distillation
Many knowledge distillation methods deploy a single-stage feature-based scheme for semantic segmentation [1, 18, 19], i.e., information is exchanged at a single network depth. Our next consideration is that cross-stage relation-based knowledge distillation is important for efficient segmentation transformers, as it enables information transfer across different network depths. Similar ideas have shown success in the past, for example, hierarchical dense prediction transformers utilize cross-stage feature maps [26, 50].
@@
We revisit Knowledge Review [2], which studies the connections across different levels of the teacher and the student CNN models. However, the Attention Based Fusion (ABF), the original feature map fusion module in Knowledge Review, dynamically aggregates feature maps via a spatial attention map, yet, overlooks the channel-wise information. We argue that the teacher models typically have more blocks for feature extraction and more channels for retaining information, but they do not differ in spatial resolution compared to student models. Therefore, we aim to bridge the channel-wise differences between the feature maps of student and teacher models while incorporating spatial information through the design of a complementary channel selection module. Our design is motivated by the following two aspects: 1) The channel-wise cues of the feature maps provide informative priors for feature extractors. Numerous previous works [73, 74], which focus on feature map fusion, seek to fuse spatial and channel information via channel attention. Furthermore, Shu et al. [1] indicate that minimizing the channel-wise Kullback-Leibler (KL) divergence of the feature- and prediction maps is more efficient than using its spatial counterpart. 2) As patch embedding distillation tracks dependencies among distant spatial positions via sequence-wise learning, feature map distillation should engage more spatially holistic knowledge in channels. It is therefore important to highlight channels from certain stages via reweighting and to account for cross-stage channel interdependencies. To achieve this, we propose the Cross Selective Fusion (CSF) module to adaptively fuse the information across stages via channel attention.
@@
Cross selective fusion. In this module, the feature maps from the current student stage and the next CSF are first resized to a uniform dimensionality and then fused via channel-wise attention. As previously demonstrated in [73], feature extractions with different kernel sizes or modalities can be fused in an adaptive way. Similarly, we propose to fuse cross-stage feature maps for a relation-based distillation. As shown in Fig. 7, we first conduct a $1{\times}1$ convolution to implement the channel-wise transformation of the input feature map $\mathbf{F}_{in}^{m}$ and conduct interpolation to implement the spatial transformation of the feature map $\mathbf{F}_{mid}^{m+1}$ from CSF of the next stage. Their height and width are the same as the feature map of the $m$-th stage $\mathbf{F}_{in}^{m}$. The resulting ${\widetilde{\mathbf{F}}^{in}}_{m}$ and ${\widetilde{\mathbf{F}}^{mid}}_{m+1}$ are element-wise summed, as depicted in Eq. (6):
@@
where $\mathbf{A},\mathbf{B}\in{\mathbb{R}}^{C\times d}$ are two learnable matrices, and $\mathbf{A}_{c},\mathbf{B}_{c}\in{\mathbb{R}}^{1\times d}$ are their $c$-th elements. $a_{c}$, $b_{c}$, and $\mathbf{F}_{mid,c}^{m}$ are the $c$-th elements of $\bm{a}$, $\bm{b}$, and $\mathbf{F}_{mid}^{m}$. $\bm{a}$ and $\bm{b}$ denote the soft attention vector for ${\widetilde{\mathbf{F}}^{in}}_{m}$ and ${\widetilde{\mathbf{F}}^{mid}}_{m+1}$, respectively. Furthermore, we perform a $3{\times}3$ convolution to transform $\mathbf{F}_{mid}^{m}$ to $\mathbf{F}_{out}^{m}$, ensuring it has the same dimension as the teacher’s feature map.
For training, we compute the hierarchical context loss (HCL) [2] between the outputs of the CSF and the teacher transformer block at each stage, as shown in Fig. 4. The feature map obtained from CSF at each stage, with shape $\left[n,c,h,w\right]$, is divided into $4$-level context information via adaptive average pooling. The height of the resulting multi-level feature maps at one stage is $\left[h,4,2,1\right]$, indicating that the shapes of the four abstract feature maps are $\left[n,c,h,w\right]$, $\left[n,c,4,4\right]$, $\left[n,c,2,2\right]$, and $\left[n,c,1,1\right]$. At each stage, the $L_{2}$ distances are utilized to distill knowledge between the context levels. The final HCL loss between the feature maps at the $m$-th stage is computed as:
$L_{fm}^{m}={\frac{1}{1+{\sum_{l}}{\frac{1}{2^{l}}}}}{\sum_{l=0}^{3}}{\frac{1}{2^{l}}}MSE(\mathbf{F}_{m,l}^{out},\mathbf{F}_{m,l}^{T}),$ (11)
where $l$ refers to the number of levels at one stage. The purpose of applying the HCL loss after every stage is to encourage knowledge distillation at different levels of abstraction.
@@ III-D Configuration
The core of our TransKD framework involves two knowledge distillation strategies: patch embedding distillation (in Sec. III-B) and feature map distillation (in Sec. III-C). To realize this design, we introduce four modules: PEA, GL-Mixer, and EA modules guide the patch embedding distillation, while the CSF feature map fusion module enhances the cross-stage relation-based feature map distillation.
Overall, CSF and PEA can be viewed as the fundamental components of TransKD. The GL-Mixer operates on the deepest representative embedding, (i.e., the final fourth stage), producing highly heterogeneous patterns [26]. It reinforces the patch embedding distillation at the final stage by harvesting global and local context information through two branches. The EA modules are placed at each stage and serve as an embedding method to provide supplementary distillation streams by using an identical embedding method to the teacher’s and the student’s, with the teacher’s number of channels. It should be taken into account that, as different transformer architectures have different embedding methods, the framework leveraging EA does not yield full plug-and-play flexibility.
We implement and examine different variants of our proposed framework. Note, that we cannot isolate and verify the efficacy of each module when all the modules are combined together, as an integration of these components means many constraints and could lead to overfitting. We, therefore, focus on three configurations of feature map distillation and patch embedding distillation provided in Tab. I, which we refer to as TransKD-Base, TransKD-GL, and TransKD-EA. The structure of these three variants is provided in Fig. 4.
Table I: Configuration of TransKD variants. FM: Feature Map. PE: Patch Embedding.
@@ IV-A Datasets and Metrics
where $k$ is the number of classes, $p_{ij}$ is the number of pixels belonging to the class $i$ and classified as class $j$. To evaluate the efficiency of TransKDs, we use MMCV [^1] MMCV: https://github.com/open-mmlab/mmcv to compute Floating point operations per second (FLOPs) and the number of parameters of our models, Additionally, we use the source code to test the Frames Per Second (FPS) of the models on an NVIDIA GeForce GTX 1080 Ti GPU. .